OpenStudy (anonymous):

help me and i will give medals! :) what is the sum of the first 10 terms of the sequence defined by an=3n-3 132 138 27 135

OpenStudy (anonymous):

someone help

OpenStudy (anonymous):

ok so what I'm GUESSING you do is replace n with numbers 1 through 10 and just solve them one by one

OpenStudy (anonymous):

for example a(1) = 3(1) - 3 a = 3-3 a=0

OpenStudy (anonymous):

that's the first one

OpenStudy (anonymous):

so would it be 27 then ?

OpenStudy (anonymous):

I honestly don't know because I would have to solve all of them, but you get the gist of what you have to do, now all you have to do is solve them :) hope I helped.